OpenStudy (anonymous):

Simplify the expression -6 × 12 + 14 − 3.

OpenStudy (anonymous):

multiply first then do the rest

OpenStudy (ddcamp):

PEMDAS: -6 x 12 = -72 -72 + 14 = -58 -58 - 3 = -61